### v1

# GPT Semantic Cache: Reducing LLM Costs and Latency via Semantic Embedding Caching

Abstract
Large Language Models (LLMs), such as GPT (Radford et al., 2019), have significantly advanced artificial intelligence by enabling sophisticated natural language understanding and generation. However, the high computational and financial costs associated with frequent API calls to these models present a substantial bottleneck, especially for applications like customer service chatbots that handle repetitive queries. In this paper, we introduce GPT Semantic Cache, a method that leverages semantic caching of query embeddings in in-memory storage (Redis). By storing embeddings of user queries, our approach efficiently identifies semantically similar questions, allowing for the retrieval of pre-generated responses without redundant API calls to the LLM. This technique reduces operational costs and improves response times, enhancing the efficiency of LLM-powered applications.
1 Introduction
Large Language Models (LLMs) like GPT have become integral to modern AI applications due to their ability to understand and generate human-like text. They are widely used in chatbots, virtual assistants, and customer support systems to interpret user queries and provide relevant responses.
Despite their capabilities, a significant challenge arises from the need to make individual API calls to the LLM for each user query. This process can be both time-consuming and costly, particularly when dealing with high volumes of similar or repetitive questions common in customer service scenarios. For example, chatbots often receive numerous inquiries that are variations of the same underlying issue, yet each is processed as a new request, incurring unnecessary computational overhead.
To address this inefficiency, we propose GPT Semantic Response, a solution that implements semantic caching of query embeddings in in-memory storage. By converting user queries into embeddings—numerical representations that capture the semantic meaning—we can store and quickly retrieve these embeddings (OpenAI, 2022). When a new query arrives, we compare its embedding to those stored in the cache. If a similar query is found, we can provide the corresponding response immediately, bypassing the need for an additional API call to the LLM.
This approach offers several advantages::
  • Reduced Latency: Users receive faster responses since the system can retrieve answers directly from the cache.
  • Cost Efficiency: Fewer API calls to the LLM result in lower operational costs.
  • Scalability: The system can handle higher volumes of queries without a proportional increase in computational resources.
2 Background
Large Language Models (LLMs), such as GPT, have significantly transformed how machines process and generate natural language text. These models are integral to applications in customer service, chatbots, and virtual assistants, where understanding and responding to queries with human-like fluency is paramount. However, using these models at scale is expensive, as each query requires a separate API call, which incurs both financial and computational costs. This inefficiency is particularly pronounced in scenarios where similar or repetitive questions are frequently asked, such as in customer support.
To address these challenges, techniques involving semantic embeddings have gained prominence (Bang, 2023). Semantic embeddings are numerical vector representations of text that encode semantic meaning, enabling efficient comparison of text data. By converting user queries into embeddings, similar queries can be identified and grouped based on their semantic closeness. Techniques such as cosine similarity are employed to measure the similarity between embeddings, allowing for the quick determination of whether a new query closely resembles a previously answered one.
3 System Architecture
The architecture of the GPT Semantic Cache system is designed to optimize the handling of repeated and semantically similar queries, reducing costs and improving response times. The system comprises several key components, each playing a crucial role in the overall workflow:
3.1 Overview
The architecture consists of three main components:
  • Embedding Generation: Converts user queries into semantic embeddings.
  • In-Memory Caching: Manages storage and retrieval of embeddings and responses using Redis.
  • Similarity Search: Identifies semantically similar queries using Approximate Nearest Neighbor (ANN) techniques.
Figure 1: The diagram illustrates the core components of the GPT Semantic Cache system, showcasing the flow of user queries through embedding generation, similarity calculation using cosine similarity (Rahutomo et al., 2012), and Approximate Nearest Neighbors (ANN) indexing. The Redis-based in-memory cache stores embeddings and corresponding responses, facilitating quick retrieval. Queries are sent to the GPT API only when no matching response is found in the cache.
3.2 Embedding Generation
The system can utilize either the OpenAI API for text embeddings or any ONNX-compatible model from platforms like Hugging Face. This flexibility allows the system to adapt to different deployment scenarios, whether using cloud-based models or self-hosted options. In the embedding generation process:
  • OpenAI API: When using OpenAI embeddings, the text is sent to the API, and the response includes the embedding vector. The vector’s length is typically 1536 for models like text-embedding-ada-002 (OpenAI, 2022).
  • Local Models: For local embeddings, the system loads ONNX-supported models and generates embeddings using libraries such as @xenova/transformers. The generated embeddings are normalized and pooled to create a consistent representation.
3.3 In-Memory Caching
The caching layer uses Redis (Labs, 2022) to store embeddings and their corresponding responses. Redis provides a high-performance, in-memory data store, ensuring quick access and efficient management of cached data. The cache is organized by:
  • Embedding Size: Different embedding models may produce vectors of varying dimensions, so the cache is partitioned based on the embedding size.
  • Time-To-Live (TTL): Cached data is assigned a TTL to prevent outdated information from persisting indefinitely. This feature ensures that the cache remains fresh and relevant, automatically expiring old entries.
3.4 Similarity Search using ANN
The system uses Approximate Nearest Neighbor (ANN) algorithms, such as Hierarchical Navigable Small World (HNSW) graphs (Malkov & Yashunin, 2018) implemented via hnswlib-node, to perform efficient similarity searches. The process involves:
  • Index Initialization: When the cache is loaded, the system initializes the ANN index with stored embeddings. If the index has not been initialized, it starts with a minimal size and dynamically grows as new data is added.
  • Embedding Search: When a new query embedding is generated, the ANN index searches for the $k$ most similar embeddings using cosine similarity. This approach ensures rapid and scalable similarity matching.
3.5 Query Handling Workflow
  1. 1.
    Cache Lookup: When a query is received, it is first converted into an embedding. The system then searches the ANN index for similar embeddings. If a match with a similarity score above a specified threshold (e.g., 0.8) is found, the cached response is returned.
  2. 2.
    Cache Miss: If no suitable match is found, the system sends the query to the LLM API (e.g., OpenAI GPT) to generate a new response. The query, embedding, and response are then stored in the cache for future use.
  3. 3.
    Embedding Storage: New embeddings and their responses are added to both the ANN index and the Redis cache. The system manages the cache size and ensures that the index can accommodate growing data by resizing when necessary.
3.6 Scalability and Model Separation
The architecture supports the separation of caches for different models or configurations, ensuring that embeddings generated by one model do not interfere with those from another. This design allows for:
  • Flexible Model Usage: The system can handle multiple embedding models simultaneously, making it adaptable to various use cases.
  • Efficient Resource Management: By using ANN and in-memory caching, the system can handle high query volumes with minimal latency, making it suitable for large-scale deployments.
3.7 Cosine Similarity and Similarity Threshold
Cosine similarity is employed to measure the similarity between query embeddings. It calculates the cosine of the angle between two vectors, providing a value between $-1$ and $1$, where $1$ indicates identical orientation. In the context of semantic embeddings, a higher cosine similarity score implies greater semantic closeness. The system uses a similarity threshold (e.g., 0.8) to decide whether a cached response is sufficiently similar to the new query to be reused.
3.8 Cache Management with Time-To-Live (TTL)
To prevent the cache from becoming stale, each cached entry is assigned a Time-To-Live (TTL) value. The TTL ensures that data is automatically removed after a certain period, maintaining the relevance of cached responses. This mechanism also helps manage the cache size, preventing it from growing indefinitely.
3.9 System Workflow Summary
The overall workflow of the GPT Semantic Cache system is as follows:
  1. 1.
    Receive Query: The user submits a query to the system.
  2. 2.
    Generate Embedding: The query is converted into an embedding using either the OpenAI API or a local ONNX model.
  3. 3.
    Search Cache: The embedding is used to search the ANN index for similar embeddings.
  4. 4.
    Determine Similarity: Cosine similarity is calculated between the query embedding and cached embeddings.
  5. 5.
    Retrieve or Generate Response:
    • Cache Hit: If a similar embedding is found above the threshold, the corresponding cached response is returned.
    • Cache Miss: If no similar embedding is found, the query is sent to the LLM API to generate a new response, which is then cached.
  6. 6.
    Update Cache and ANN Index: The new embedding and response are stored in Redis and added to the ANN index for future queries.
3.10 Advantages of the Architecture
The GPT Semantic Cache system offers several benefits:
  • Reduced Latency: By serving responses from the cache when possible, the system provides faster response times to users.
  • Cost Savings: Reducing the number of API calls to the LLM lowers operational costs significantly.
  • Scalability: The use of ANN and in-memory caching allows the system to handle increasing volumes of queries efficiently.
  • Flexibility: Support for multiple embedding models and configurations makes the system adaptable to various deployment needs.
3.11 Potential Extensions
Future enhancements to the system could include:
  • Dynamic Threshold Adjustment: Implementing mechanisms to adjust the similarity threshold based on system performance or user feedback.
  • Distributed Caching: Expanding the caching layer to a distributed setup for higher availability and fault tolerance.
  • Advanced Embedding Models: Incorporating newer or domain-specific embedding models to improve semantic understanding.
4 Testing Methodology
To evaluate the effectiveness of our GPT Semantic Cache system, we conducted a series of experiments designed to simulate real-world user interactions with AI-based services. The primary objectives were to assess the system’s ability to efficiently handle repeated or semantically similar queries, validate the accuracy of cached responses, and quantify improvements in both response time and operational cost.
4.1 Dataset Preparation and Cache Population
We constructed a dataset of 8,000 question-answer pairs across four categories: basic Python programming, customer service inquiries, technical support questions, and general knowledge. These categories were selected to reflect the diverse range of queries typically encountered in real-world applications. Each question was converted into a semantic embedding using all-MiniLM-L6-v2 (Reimers & Gurevych, 2020), and both the embeddings and their corresponding responses were stored in Redis. An Approximate Nearest Neighbor (ANN) index was also updated to include these embeddings for efficient similarity searches.
4.2 Test Query Generation and Execution
To evaluate the system’s performance, we generated an additional 2,000 test queries, with 500 queries per category. These test queries were designed to have a cosine similarity of at least 0.8 with one or more cached questions, ensuring that they were semantically similar. Each query was submitted to the system, where it was converted into an embedding and compared against cached embeddings using the ANN index. If a similar embedding was found (similarity ¿= 0.8), the corresponding cached response was returned (cache hit). Otherwise, the query was forwarded to the LLM for a new response (cache miss), which was subsequently cached.
4.3 Validation of Cache Hits
To verify the appropriateness of cached responses, we employed GPT-4o Mini to evaluate whether the retrieved cached response was valid for the test query. For each cache hit, both the test query and the original cached question were provided to the model, which returned a binary verdict indicating whether the queries were semantically similar and whether the cached response was accurate. This validation step enabled us to quantify the accuracy of the caching mechanism.
5 Results
The evaluation of GPT Semantic Cache highlights its significant impact on reducing API calls and improving response times. This section provides a detailed analysis of system performance, comparing the traditional method to our caching approach.
5.1 Reduction in API Calls
GPT Semantic Cache reduces the number of API calls by serving responses from the cache for semantically similar queries. As shown in Figure 1, the traditional method results in 100% API calls across all categories, whereas GPT Semantic Cache reduces API dependency significantly. The reduction rates for each category are as follows:
  • Basics of Python Programming: 67% cache hit rate, reducing API calls to 33%.
  • Technical Support Related to Network: 67% cache hit rate, reducing API calls to 33%.
  • Questions Related to Order and Shipping: 68.8% cache hit rate, reducing API calls to 31.2%.
  • Customer Shopping QA: 61.6% cache hit rate, reducing API calls to 38.4%.
Figure 2: Reduction in API Calls: Traditional Method vs GPT Semantic Cache.
5.2 Cache Hit and Accuracy Analysis
The cache hit rates indicate the system’s efficiency in recognizing semantically similar queries, while the positive hit rates highlight the accuracy of responses retrieved from the cache. As presented in Figure 2, the system achieved high positive hit rates across all categories, ranging from 92.5% to 97.3%, confirming the reliability of cached responses.
5.3 Performance Metrics
Figure 3: Cache Hit Percentage per Category
Table 1: Cache Hit per 500 queries in each category and the no of positive hits
Category Cache Hit Positive Hits
Basics of Python Programming 335 310
Technical Support Related to Network 335 326
Questions Related to Order and Shipping 344 331
Customer Shopping QA 308 298
6 Discussion
The results clearly illustrate the advantages of GPT Semantic Cache in reducing API dependency while maintaining high response accuracy. By achieving cache hit rates up to 68.8% and positive hit rates exceeding 97%, the system not only boosts efficiency but also ensures the reliability of responses retrieved from the cache. This high hit rate means that a significant portion of queries can be served directly from the cache, bypassing the need for costly and time-consuming API calls to the language model. This reduction in API calls translates directly into lower operational costs and faster response times, which are critical in high-demand applications such as customer service and technical support. The consistent accuracy of cached responses across categories also underscores the system’s reliability, making it suitable for deployment in environments where response quality and relevance are essential.
In addition to cost efficiency, GPT Semantic Cache demonstrates substantial scalability and adaptability. By using Approximate Nearest Neighbor (ANN) algorithms and in-memory caching with Redis, the system can handle large volumes of queries without a proportional increase in computational demands, supporting stable performance even under fluctuating workloads. The system’s use of a Time-To-Live (TTL) mechanism also ensures that cached data remains fresh, further enhancing the relevance and reliability of responses in dynamic environments. Together, these features highlight GPT Semantic Cache as a viable, cost-effective, and scalable solution for large-scale AI applications requiring quick, accurate, and economical query handling.
Appendix A Appendix
The Package, Dataset and Code used to write this paper